a snowboarder in a race is travelling at 10 m/s as he crosses the finishing line. he then decelerates uniformly until coming to a rest over a distance of 20 m. What is his acceleration as he pulls up
1
Expert's answer
2018-04-24T09:39:08-0400
v_f^2-v_i^2=2as
0^2-〖10〗^2=2a(20)
a=-100/40=-2.5 m/s^2 .